Cary is 7 years older than Dan. In 3 years the sum of their ages will be 81 . Find the age of each man now
Answer by richwmiller(17219)

Dan=(81-6+7)/2
Dan=(82)/2
Dan=41
Cary=41+7
Dan=48
Dan=41 Cary= 48